reuben has scores 70,75,83,and 80 on four spanish quizzes.what score mush he get on next quiz average of atleast 80?
Answer by Alan3354(69443)

Here's a similar problem.
=============================
John received three grades of 85, 92, and 100 on his first three tests. What must he get on his fourth test to get a 90 average?
------------
He has 85+92+100 = 277 "points"
4*90 = 360
360 - 277 = 83